OpenStudy (anonymous):

How to solve this log x - log 12 = 3

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \log(x)-\log(b)=\log(x/b) } \]

OpenStudy (solomonzelman):

\[\color{blue}{ \log(x)-\log(12)=\log(x/12) } \]\[So,~~~~~~~~~~\color{blue}{ \log(x/12)=3 } \]\[\color{blue}{ \log(x/12)=(\log_33) \times 3 } \]\[\color{blue}{ \log(x/12)=3(\log_33) } \]\[\color{blue}{ \log(x/12)=(\log_33^3) } \]\[\color{blue}{ \log(x/12)=\log_327 } \]\[\color{blue}{ \log(x/12)=\frac{\log_{10}27}{\log_{10}3} } \]can you do it from here?